• install and test new limit switches on ARGOS calibration source swing arms (SX & DX)
  • removal of temporary diagnostic lines on SX ARGOS laser boxes, racks and LGSW (closed IT 6060)
  • installation of the analog summing amplifier on SX&DX, to improve vibration compensation
  • characterization of the installed piezo amplifier
  • LM1 heater firmware update
  • removal of the temporary accelerometer cables on SX
  • installation of snap ferrites on the pnCCD sequencer cables to improve EMC situation
  • arrival of additional ARGOS technical staff today (Marco, Jose and Mathias)
  • arrival of Chris Biddick for software support
Topic revision: r1 - 28 Apr 2017, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback